### v5

# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

Abstract
Efficiently learning and executing long-horizon mobile manipulation (MoMa) tasks is crucial for advancing robotics in household and workplace settings. However, current MoMa models are data-inefficient, underscoring the need for improved models that require realistic-sized benchmarks to evaluate their efficiency, which do not exist. To address this, we introduce the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities), which evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. The benchmark includes 571 human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We benchmark several models, including learning-based models and a neuro-symbolic modular approach combining foundation models with task and motion planning. Learning-based models show suboptimal success rates, even when leveraging pretrained weights, underscoring significant data inefficiencies. However, the neuro-symbolic approach performs significantly better while being more data efficient. Findings highlight the need for more data-efficient learning-based MoMa approaches. addresses this gap by serving as a key benchmark for evaluating the data efficiency of those future models in handling household robotics tasks.
[…]
figureOverview. LAMBDA () is a benchmark comprised of 571 language-conditioned long-horizon (room-to-room and floor-to-floor navigation) mobile manipulation (pick-and-place) tasks paired with human-collected expert demonstrations from both simulated and real-world environments. The illustrated examples in both the simulated and real-world environments span multiple rooms, which is a skill expected of daily household and workspace robots. We use a quadruped robot for its ability to recover from falls and to navigate complex terrain such as stairs, which is essential for floor-to-floor tasks.
I INTRODUCTION
Improving efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6]. Collecting such large-scale datasets is resource-intensive and requires substantial time, manpower, and infrastructure.This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse task requirements, and the need for generalization in home and workplace settings. Addressing these challenges demands the development of more data-efficient approaches for MoMa, supported by benchmarks that can guide and evaluate this progress effectively.
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, for long-horizon, language-conditioned MoMa tasks, leaving a significant gap in evaluating and advancing progress in this area. For instance, most lack language instruction annotations for the tasks [7, 8], even though natural language provides an intuitive and accessible interface for specifying robot objectives, especially for household and workplace users. Also, most benchmarks do not support long-horizon tasks [4, 9] requiring navigation across rooms and floors, which are essential for realistic household scenarios. Furthermore, while some benchmarks support simulated environments, they often lack real-world data [10, 11], which are crucial for ensuring models’ effectiveness in practical deployments. Some current benchmark datasets also rely on planner-generated trajectories [12, 13], which lack the natural variability and expertise of human-collected demonstrations. Finally, many benchmarks lack quadruped data [14, 15], which are crucial for terrains like stairs for floor-to-floor navigation, and maintaining stability during tasks.
To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. The dataset consists of 571 high-quality trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in provide natural variability and expertise, better capturing real-world robot operations. Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Its simulated environments enable diverse testing scenarios and object variability, ensuring flexibility and standardization, while real-world data ensures models are prepared for practical deployment. Finally, the inclusion of armed quadrupeds enables evaluation on complex terrains, such as stairs, highlighting their stability and adaptability in challenging conditions.
We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the dataset’s ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future. These results demonstrate ’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved ones for indoor robotics.
Our contributions are summarized as follows:
benchmark: We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
High-quality dataset: ’s tasks are paired with 571 human-collected trajectories, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
Comprehensive evaluation: We benchmark several models, including BC and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
| Benchmark | Natural Language | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
| SafeAgentBench | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| HumanoidBench | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ |
| Mobile ALOHA | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| BEHAVIOR-1K | ✗ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ |
| ManiSkill-HAB | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| InfiniteWorld | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| RoboSuite | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✓ |
| RoboCasa | Some | ✗ | ✓ | Some | ✓ | ✓ | ✓ |
| BRMData | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| DeMoBot | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MoMaRT | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
| M3Bench | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| BiGym | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
| RT-1 | ✓ | ✗ | ✓ | ✓ | ✓ | ✓ | ✗ |
II Related Work
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [7, 8, 16, 13, 17, 18, 19, 15, 9], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications.
Moreover, almost all benchmarks [7, 8, 13, 17, 14, 18, 19, 15, 12, 9, 4], do not support long-horizon tasks involving navigation across rooms and floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. While most benchmarks provide simulated environments, Mobile ALOHA [8] and BRMData [18] lack simulation data entirely. Simulated environments offer diverse scenes and objects for flexible testing across varied scenarios and provide a standardized, portable testbed that can be easily used by others due to their fixed conditions. On the other hand, most lack real-world data [10, 7, 13, 11, 17, 15, 12, 9], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
Most benchmarks provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [10], HumanoidBench [7], BEHAVIOR-1K [16], and InfiniteWorld [11], lack this attribute, limiting their utility for model training and evaluation. M3Bench [12] and ManiSkill-HAB [13] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
Finally, most benchmarks from Table I, with the exceptions of RoboCasa [14], RoboSuite [17], and DeMoBot [19], do not feature armed quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in . Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
By addressing these limitations, the benchmark provides a comprehensive benchmark to test the data efficiency of MoMa models. It integrates language-conditioned tasks, long-horizon trajectories spanning multiple rooms and floors, and both simulated and real-world data. Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed quadruped data makes it uniquely suited for advancing robotics in complex real-world settings.
III Benchmark
We present LAMBDA (), a benchmark designed to evaluate model data efficiency on long-horizon MoMa tasks. features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings. The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
III-A Challenges of Tasks
Data Efficiency in MoMa. is designed to test the data efficiency of learning-based models by offering a dataset of moderate size, deliberately curated to strike a balance between realism and practicality. While recent data-efficient tabletop manipulation models [1, 2, 3] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa. MoMa encompasses not only higher-dimensional observation, state, and action spaces but also significantly larger and more varied task spaces. This deliberate dataset size ensures sufficient diversity for learning while maintaining a manageable scale that avoids the excessive computational and collection costs of unrealistically large datasets.
Complex Spaces. The inclusion of MoMa tasks introduces significant complexity due to the interplay of action, observation, state, and task spaces. Unlike fixed single-arm setups, MoMa requires navigating and interacting within dynamic three-dimensional environments, often spanning multiple rooms or floors. This necessitates the seamless integration of high-level planning and low-level motion control, creating challenges for robotic models to operate effectively. For example, consider a household task where the agent must navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this straightforward pick-and-place task, the complexity of the action, state, and observation spaces introduces substantial variability. The agent might approach the table from different directions, choose from various grasping points on the plate, or follow alternative routes to the destination. This multi-modal nature means that the same task can be solved in multiple ways, further increasing the demands on models to generalize across diverse strategies and environmental configurations. Moreover, the complexity of these integrated spaces enables a broader range of tasks, from transporting objects across floors to rearranging items in dynamic environments. For our benchmark, we focus on pick-and-place tasks, as they represent a significant portion of daily activities and provide a foundational challenge for MoMa models to address.
Long-Horizon Tasks. A key challenge in lies in its dedicated emphasis on long-horizon tasks, where navigation spans multiple rooms and floors in indoor environments such as households and workplaces. These tasks inherently demand long-term planning, often requiring hierarchical strategies that integrate high-level task objectives with low-level motion execution. The agent must account for spatial relationships, sequential task dependencies, and the compounding effects of errors over time. Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in BC. For instance, navigating to a distant room to retrieve an object and returning it to another location involves traversing obstacles and staircases while adhering to task objectives and maintaining situational awareness. Such tasks are representative of real-world scenarios where robots are expected to perform complex, multi-step instructions over extended periods.
Realistic Task Distribution. To ensure the tasks in reflect practical needs, we used crowdsourcing to gather a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. This method avoided templated task specifications by capturing natural, human-generated instructions. The resulting dataset encompasses a variety of pick-and-place MoMa tasks, aligning with real-world expectations and exposing models to inherently varied and context-dependent scenarios.
III-B Simulated Dataset
The dataset for our simulated tasks consists of 521 trajectories collected across 20 rooms in five distinct environments, designed to ensure diversity in both object types and room layouts. We utilized the AI2THOR [20] simulator and specifically employed RoboTHOR [21] environments, which feature multiple interconnected rooms suitable for long-horizon tasks.
For robot control, we used ManipulaTHOR [22], which includes an arm capable of fine manipulation and provides low-level pose data. Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions taken during the tasks.
To generate realistic and diverse instructions, we leveraged Prolific111prolific.com for crowdsourcing to gather 521 natural language commands from 41 participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. Using these commands, we teleoperated the robot in the simulator with a custom keyboard-based teleoperation system we developed, ensuring accuracy and consistency throughout the data collection process. After teleoperation, each trajectory was replay-verified to confirm correctness and eliminate errors, ensuring high-quality data integrity.
III-C Real-World Dataset
The dataset for our real-world tasks comprises 50 trajectories collected across 10 rooms in three distinct environments. These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities. The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building. These spaces include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to enhance the generalizability of trained models.
To handle complex terrain and ensure stability, we selected the Spot [23] quadruped mobile manipulator, equipped with an arm for pick-and-place tasks. The quadruped’s design enables it to navigate stairs and maintain balance during task execution, making it ideal for the challenging settings of our chosen environments.
We teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on 50 natural language commands provided by seven participants. The commands, distributed as 20, 15, and 15 across the three environments, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings. Task trajectories were collected at 3 Hz to ensure detailed and realistic data.
IV Evaluation
IV-A Baseline Models
To evaluate the and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models. Given the growing popularity of imitation learning (IL), particularly BC, we benchmark two state-of-the-art (SOTA) BC models. These models provide insight into how data-efficient current BC algorithms are by being tested on , validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success. Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). This system, evaluated zero-shot without requiring training, investigates whether non-BC methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies. Finally, we assess the performance of a random agent, wherein actions are uniformly sampled to establish a lower-bound baseline.
RT-1. Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a Transformer architecture to process natural language instructions and RGB image inputs into discretized actions. Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robot’s capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation to achieve precise control.
MotionGlot-MoMa MG [24] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [25]. Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem. We adapt MotionGlot (MG) for MoMa tasks by leveraging ’s demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa. MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions. This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [26]. This unified vocabulary allows for action generation similar to text generation. Finally, the instruction-tuning template from MG-MoMa is used for text command-based action generation.
LIMP. Language Instruction grounding for Motion Planning (LIMP) [27] is a neuro-symbolic system for pick-and-place MoMa that integrates Linear Temporal Logic (LTL), foundation models, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating natural language instructions into LTL formulas enriched with Composable Referent Descriptors, LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
| Scene 1 | Scene 2 | Scene 3 | Scene 4 | Scene 5 | Task Gen | |||||||||||||
| Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | |
| Demos | 330 | 83 | 108 | 340 | 86 | 95 | 337 | 85 | 99 | 324 | 82 | 115 | 333 | 84 | 104 | 416 | 51 | 54 |
| Scenes | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 5 | 5 | 5 |
IV-B Experiments.
Data Split & Generalization. To evaluate generalization capabilities, we conduct experiments using carefully designed data splits for scene generalization and task generalization on the simulation dataset, highlighted in Table II.
For scene generalization, we employ a 5-fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, 80% of the trajectories from each of the four remaining scenes are included in the training set, and 20% are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene. This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of scene generalization across the entire dataset.
For task generalization, the data split is consistent across all five scenes. The training set comprises 80% of the trajectories from each scene, the validation set includes 10%, and the test set includes the remaining 10%. Unlike scene generalization, the held-out test set in this experiment evaluates the model’s ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for task generalization on the test set.
Training. To benchmark the data efficiency of modern BC methods, we train the RT-1 and MotionGlot models from scratch without any pretrained parameters on the simulation tasks. This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks. Both models are trained using the scene generalization and task generalization experiments, ensuring their capabilities are evaluated under varying conditions.
Fine-tuning. To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the simulation tasks. Empirical analysis showed that freezing some model layers led to better performance than not freezing any layers, so this approach was used for the final experiment. Using this setup, we conducted the task generalization experiment to measure how much pretrained MoMa knowledge impacts RT-1’s ability to adapt efficiently to new MoMa tasks.
V Discussion
V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively. None of the individual scenes achieved success rates exceeding , indicating significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for long-horizon MoMa tasks. Both models outperform the random agent, demonstrating their ability to learn from the data and satisfy the minimum performance criterion.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robot’s initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 1 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
Interestingly, for MG-MoMa, Scenes 1 and 3 were the exact opposite of RT-1, with those being the lowest performers. It is unclear why that is the case, but it could be that MG-MoMa fails to recognize nearby objects. Similarly, Mg-MoMa’s best performers are RT-1’s worst, Scenes 2 and 5. Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that Motionglot was not originally intended for mobile manipulation tasks.
Task Generalization. The success rates, presented in Table IV, for the Task Generalization experiment for RT-1 and Mg-MoMa, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment. This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments. Additionally, while it is unsurprising that RT-1 outperformed MG-MoMa on unseen demonstrations, it is notable that the two models performed comparably on seen demonstrations, and the margin of RT-1’s outperformance was relatively small. This is remarkable, particularly given that MotionGlot was originally designed for cross-embodied motion generation rather than specifically for mobile manipulation.
We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks. Surprisingly, RT-1 achieved a marginally higher success rate on unseen tasks () compared to seen tasks (). This indicates that the model’s training did not overfit to the training data, which is a positive outcome. One plausible explanation is that the sampler may have inadvertently selected a disproportionate number of especially challenging tasks for the seen evaluation set, rendering them overly difficult for the model despite having been part of the training data.
Task Generalization with Fine-Tuning. Interestingly, the unseen task performance was marginally lower compared to the fully trained model’s performance on unseen tasks, as seen in Table IV. This result suggests the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world.
LIMP. The results presented thus far indicate that learning-based models, face significant challenges in addressing the complexities of long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, assessing 45 tasks, including a mix of tasks originally specified for both simulated and real-world environments. Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks. For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. This enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training, making it relatively more data efficient.
One key factor contributing to LIMP’s success is its use of pre-generated maps of the environments. They provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless autoregressive action decoding used in BC models like RT-1. This robustness is critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive prior knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables LIMP to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training. While LIMP demonstrates considerable promise, it still falls short of perfect task execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. Overall, LIMP represents a promising direction for future system development, blending the strengths of learning-based approaches with those of classical planning.
V-B Ablations
Impact of Dataset Size. The poor performance observed with learning-based methods is expected to stem from the limited, realistically sized dataset of the benchmark, but this needs to be further examined. Will simply scaling the data improve the performance? Will these models effectively and efficiently use more data to learn better? To address this, we investigate how performance scales when progressively larger subsets of the dataset (, , , and ) are used for training. This evaluates whether these models can effectively utilize limited data or require substantially more, directly testing their data efficiency and scalability for long-horizon MoMa tasks.
The ablation results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset. We attribute part of the variation to the inherent complexity of these long-horizon MoMa tasks and the variability in the sampled subsets, where smaller subsets can incidentally include a higher proportion of comparatively easier scenarios. Overall, this suggests that simply providing more samples does not guarantee better learning outcomes for these tasks, and it is likely that limited data is not the primary cause of the poor performance. Moreover, since the linear increase in data size did not result in proportional improvement, it indicates that the model is unable to efficiently learn from the additional data provided, deeming it data inefficient. This raises questions about deeper limitations, such as architectural constraints, warranting further investigation.
| Scene Gen | ||||||
| Model | Scene 1 | Scene 2 | Scene 3 | Scene 4 | Scene 5 | Avg |
| RT-1 | ||||||
| MG-MoMa | ||||||
| Random | ||||||
| Task Gen | Task Gen FT | |||
| Model | Seen | Unseen | Seen | Unseen |
| RT-1 | ||||
| MG-MoMa | – | – | ||
| LIMP | – | – | – | |
| Random | – | – | ||
Impact of Architecture. To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture. While MotionGlot was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also relies on a Transformer architecture. To explore a more distinct alternative, we developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks. The LSTM variant is evaluated under the task generalization setup using subset experiments with , , , and of the dataset, enabling us to analyze how its performance scales with increasing data. By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
As shown in Table V, the LSTM variant exhibits no upward trend when moving from to of the data, indicating it is also data inefficient. To compare the trends between the two architectures, we analyze their average rates of change. Both exhibit similarly non-positive trends, indicating that the data inefficiency is unlikely to primarily stem from the choice of architecture and seems to span across different architectures, since they are not efficiently making use of more data. Furthermore, since RT-1’s performance is somewhat better than the LSTM variant’s, and its RoC is slightly higher, it indicates marginally better data efficiency, further supporting the argument that architecture is unlikely the primary cause of RT-1’s data inefficiency observed in the main experiments.
Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are simply inherently too complex and challenging for learning-based methods. Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this work and is left for the community to explore in future work.
VI Conclusion
We introduce LAMBDA (), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments. Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
To evaluate the benchmark, we tested multiple models to assess their performance and data efficiency on these complex tasks. Learning-based models, RT-1 and MotionGlot, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to ’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by ’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
Limitations and Future Work. While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation. Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models. Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency. Evaluating their performance on ’s sparse rewards would offer additional insights into their suitability for long-horizon MoMa tasks and broaden the scope of the benchmark.
| Model | Avg RoC | ||||
| RT-1 | |||||
| LSTM |

### v6

@@
 Abstract
-Efficiently learning and executing long-horizon mobile manipulation (MoMa) tasks is crucial for advancing robotics in household and workplace settings.
-However, current MoMa models are data-inefficient, underscoring the need for improved models that require realistic-sized benchmarks to evaluate their efficiency, which do not exist.
-To address this, we introduce the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities), which evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection.
-The benchmark includes 571 human-collected demonstrations that provide realism and diversity in simulated and real-world settings.
+Learning to execute long-horizon mobile manipulation tasks is crucial for advancing robotics in household and workplace settings.
+However, current approaches are typically data-inefficient, underscoring the need for improved models that require realistically sized benchmarks to evaluate their efficiency.
+To address this, we introduce the LAMBDA () benchmark111lambdabenchmark.github.ioââLong-horizon Actions for Mobile-manipulation Benchmarking of Directed Activitiesââwhich evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection.
+Our benchmark includes human-collected demonstrations that provide realism and diversity in simulated and real-world settings.
 Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation.
-We benchmark several models, including learning-based models and a neuro-symbolic modular approach combining foundation models with task and motion planning.
-Learning-based models show suboptimal success rates, even when leveraging pretrained weights, underscoring significant data inefficiencies.
-However, the neuro-symbolic approach performs significantly better while being more data efficient.
-Findings highlight the need for more data-efficient learning-based MoMa approaches.
-addresses this gap by serving as a key benchmark for evaluating the data efficiency of those future models in handling household robotics tasks.
-[…]
-figureOverview.
-LAMBDA () is a benchmark comprised of 571 language-conditioned long-horizon (room-to-room and floor-to-floor navigation) mobile manipulation (pick-and-place) tasks paired with human-collected expert demonstrations from both simulated and real-world environments.
-The illustrated examples in both the simulated and real-world environments span multiple rooms, which is a skill expected of daily household and workspace robots.
-We use a quadruped robot for its ability to recover from falls and to navigate complex terrain such as stairs, which is essential for floor-to-floor tasks.
+We leverage LAMBDA to benchmark current end-to-end learning methods and a modular neuro-symbolic approaches that combines foundation models with task and motion planning.
+We find that end-to-end methodsâeven when pretrainedâyield lower success rates, while neuro-symbolic methods perform significantly better and require less data.
 I INTRODUCTION
-Improving efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications.
+Improving data efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications.
 These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions.
-Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6].
-Collecting such large-scale datasets is resource-intensive and requires substantial time, manpower, and infrastructure.This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse task requirements, and the need for generalization in home and workplace settings.
-Addressing these challenges demands the development of more data-efficient approaches for MoMa, supported by benchmarks that can guide and evaluate this progress effectively.
-Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, for long-horizon, language-conditioned MoMa tasks, leaving a significant gap in evaluating and advancing progress in this area.
-For instance, most lack language instruction annotations for the tasks [7, 8], even though natural language provides an intuitive and accessible interface for specifying robot objectives, especially for household and workplace users.
-Also, most benchmarks do not support long-horizon tasks [4, 9] requiring navigation across rooms and floors, which are essential for realistic household scenarios.
-Furthermore, while some benchmarks support simulated environments, they often lack real-world data [10, 11], which are crucial for ensuring models’ effectiveness in practical deployments.
-Some current benchmark datasets also rely on planner-generated trajectories [12, 13], which lack the natural variability and expertise of human-collected demonstrations.
-Finally, many benchmarks lack quadruped data [14, 15], which are crucial for terrains like stairs for floor-to-floor navigation, and maintaining stability during tasks.
+Significant strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3], however existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6].
+Collecting such large-scale datasets is resource-intensive, requiring substantial time, labor, and infrastructure.
+This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse tasks, and the need for generalization across environments.
+Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, leaving a significant gap in evaluating progress in this area.
+A few works  [7, 8, 9, 10, 4, 11] pair task demonstrations with language instruction annotations and support long-horizon tasks [7, 12, 8, 11], providing an intuitive and accessible interface for specifying robot objectives.
+Even fewer [9, 11] provide natural language instructions in an open-ended, flexible format rather than relying on fixed templates.
+Many benchmarks lack real-world data [13, 12, 9, 14, 15, 4] or rely on planner-generated trajectories [16, 10, 17], limiting their practicality.
+Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide quadruped data.
 To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities).
 evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks.
-The dataset consists of 571 high-quality trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability.
+Our dataset consists of trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability.
 Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments.
@@
 Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios.
-Its simulated environments enable diverse testing scenarios and object variability, ensuring flexibility and standardization, while real-world data ensures models are prepared for practical deployment.
-Finally, the inclusion of armed quadrupeds enables evaluation on complex terrains, such as stairs, highlighting their stability and adaptability in challenging conditions.
+Simulated environments enable diverse testing scenarios and object variability, while real-world data ensures models are prepared for practical deployment.
+We collected this dataset with a quadruped robot [18] due to its optimal form-factor suitable for navigating complex terrain.
 We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks.
-Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the dataset’s ability to support learning.
-A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future.
-These results demonstrate ’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved ones for indoor robotics.
-Our contributions are summarized as follows:
-benchmark: We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
-High-quality dataset: ’s tasks are paired with 571 human-collected trajectories, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
-Comprehensive evaluation: We benchmark several models, including BC and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
-| Benchmark | Natural Language | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
-| SafeAgentBench | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
-| HumanoidBench | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ |
-| Mobile ALOHA | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
-| BEHAVIOR-1K | ✗ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ |
-| ManiSkill-HAB | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
-| InfiniteWorld | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
-| RoboSuite | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✓ |
-| RoboCasa | Some | ✗ | ✓ | Some | ✓ | ✓ | ✓ |
-| BRMData | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
-| DeMoBot | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
-| MoMaRT | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
-| M3Bench | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
-| BiGym | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
-| RT-1 | ✓ | ✗ | ✓ | ✓ | ✓ | ✓ | ✗ |
+Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the datasetâs ability to support learning.
+A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
+Overall our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
+(2) High-quality dataset:âs tasks are paired with human-collected robot demonstrations, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
+(3) Comprehensive evaluation:We benchmark several models, including BC and modular neuro-symbolic approaches, providing a baseline for future model development for MoMa tasks.
+| Benchmark | NL | FF NL | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
+| SafeAgentBench [7] | â | â | â | â | â | â | â | â |
+| HumanoidBench [20] | â | â | â | â | â | â | â | â |
+| Mobile ALOHAÂ [13] | â | â | â | â | â | â | â | â |
+| BEHAVIOR-1KÂ [12] | â | â | â | â | â | â | â | â |
+| ManiSkill-HABÂ [16] | â | â | â | â | â | â | â | â |
+| InfiniteWorld [8] | â | â | â | â | â | â | â | â |
+| M2Diffuser [17] | â | â | â | â | â | â | â | â |
+| RoboSuite [19] | â | â | â | â | â | â | â | â |
+| BRMData [14] | â | â | â | â | â | â | â | â |
+| RoboCasa [9] | Some | â | â | â | Some | â | â | â |
+| DeMoBot [15] | â | â | â | â | â | â | â | â |
+| MoMaRTÂ [21] | â | â | â | â | â | â | â | â |
+| M3Bench [10] | â | â | â | â | â | â | â | â |
+| PARTNRÂ [11] | â | â | â | â | â | â | â | â |
+| BiGym [22] | â | â | â | â | â | â | â | â |
+| RT-1Â [4] | â | â | â | â | â | â | â | â |
+| â | â | â | â | â | â | â | â |
 II Related Work
@@
 Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications.
-Moreover, almost all benchmarks [7, 8, 13, 17, 14, 18, 19, 15, 12, 9, 4], do not support long-horizon tasks involving navigation across rooms and floors.
+Even fewer benchmarks, such as RoboCasa [9] and PARTNR [11], offer natural language instructions in a free-form, unconstrained format rather than predefined templates.
+This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in-the-wild.
+Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17], do not support long-horizon tasks involving navigation across rooms and floors.
 These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations.
 Without this capability, benchmarks fail to reflect the complexity of real-world scenarios.
-While most benchmarks provide simulated environments, Mobile ALOHA [8] and BRMData [18] lack simulation data entirely.
-Simulated environments offer diverse scenes and objects for flexible testing across varied scenarios and provide a standardized, portable testbed that can be easily used by others due to their fixed conditions.
-On the other hand, most lack real-world data [10, 7, 13, 11, 17, 15, 12, 9], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
-Most benchmarks provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models.
-However, a few, SafeAgentBench [10], HumanoidBench [7], BEHAVIOR-1K [16], and InfiniteWorld [11], lack this attribute, limiting their utility for model training and evaluation.
-M3Bench [12] and ManiSkill-HAB [13] depend on planner-generated trajectories.
+As shown in Table I, Mobile ALOHA [13] and BRMData [14] lack simulation data entirely.
+Simulated environments offer diverse scenes and objects, making it easier to test models across a wide range of scenarios.
+Their standardized and controlled conditions create a portable testbed that others can readily use, ensuring consistency and reproducibility in evaluations.
+On the other hand, most lack real-world data [7, 20, 16, 8, 19, 21, 10, 22, 11, 17], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
+The majority of benchmarks in Table I provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models.
+However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11] lack this attribute, limiting their utility for model training and evaluation.
+M3Bench [10], M2Diffuser [17], and ManiSkill-HAB [16] depend on planner-generated trajectories.
 These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
-Finally, most benchmarks from Table I, with the exceptions of RoboCasa [14], RoboSuite [17], and DeMoBot [19], do not feature armed quadruped data, which is essential for tasks requiring stability and adaptability.
+Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed quadruped data, which is essential for tasks requiring stability and adaptability.
 Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in .
@@
 III Benchmark
-We present LAMBDA (), a benchmark designed to evaluate model data efficiency on long-horizon MoMa tasks.
+The LAMBDA () benchmark is designed to evaluate model data efficiency on long-horizon MoMa tasks.
 features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings.
@@
 Data Efficiency in MoMa.
-is designed to test the data efficiency of learning-based models by offering a dataset of moderate size, deliberately curated to strike a balance between realism and practicality.
-While recent data-efficient tabletop manipulation models [1, 2, 3] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa.
-MoMa encompasses not only higher-dimensional observation, state, and action spaces but also significantly larger and more varied task spaces.
-This deliberate dataset size ensures sufficient diversity for learning while maintaining a manageable scale that avoids the excessive computational and collection costs of unrealistically large datasets.
-Complex Spaces.
-The inclusion of MoMa tasks introduces significant complexity due to the interplay of action, observation, state, and task spaces.
-Unlike fixed single-arm setups, MoMa requires navigating and interacting within dynamic three-dimensional environments, often spanning multiple rooms or floors.
-This necessitates the seamless integration of high-level planning and low-level motion control, creating challenges for robotic models to operate effectively.
-For example, consider a household task where the agent must navigate to a dining table, pick up a plate, and transport it to a countertop in another room.
-Even in this straightforward pick-and-place task, the complexity of the action, state, and observation spaces introduces substantial variability.
-The agent might approach the table from different directions, choose from various grasping points on the plate, or follow alternative routes to the destination.
-This multi-modal nature means that the same task can be solved in multiple ways, further increasing the demands on models to generalize across diverse strategies and environmental configurations.
-Moreover, the complexity of these integrated spaces enables a broader range of tasks, from transporting objects across floors to rearranging items in dynamic environments.
-For our benchmark, we focus on pick-and-place tasks, as they represent a significant portion of daily activities and provide a foundational challenge for MoMa models to address.
-Long-Horizon Tasks.
-A key challenge in lies in its dedicated emphasis on long-horizon tasks, where navigation spans multiple rooms and floors in indoor environments such as households and workplaces.
-These tasks inherently demand long-term planning, often requiring hierarchical strategies that integrate high-level task objectives with low-level motion execution.
-The agent must account for spatial relationships, sequential task dependencies, and the compounding effects of errors over time.
-Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in BC.
-For instance, navigating to a distant room to retrieve an object and returning it to another location involves traversing obstacles and staircases while adhering to task objectives and maintaining situational awareness.
-Such tasks are representative of real-world scenarios where robots are expected to perform complex, multi-step instructions over extended periods.
-Realistic Task Distribution.
-To ensure the tasks in reflect practical needs, we used crowdsourcing to gather a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings.
-This method avoided templated task specifications by capturing natural, human-generated instructions.
-The resulting dataset encompasses a variety of pick-and-place MoMa tasks, aligning with real-world expectations and exposing models to inherently varied and context-dependent scenarios.
+We define data efficiency as the ability to achieve relatively strong performance with minimal robot trajectory demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every time step.
+Non-end-to-end systems are still considered within our definition of data efficiency.
+For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstration trajectories, are not excluded on this basis.
+Unlike very large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
+While previous tabletop manipulation methods depended on thousands of demonstrations [23], recent models [1, 2, 3] are more data efficient, requiring only tens of demonstrations.
+In a similar direction, recognizing the increased complexity of MoMa, we provide hundreds of demonstrations for model development and benchmarking.
+This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient, avoiding the excessive resource costs of unrealistically large datasets.
+Long Horizon Tasks in Complex Spaces.
+MoMa tasks introduce significant complexity by involving larger and more intricate action, observation, state, and task spaces compared to tabletop manipulation.
+Operating in dynamic three-dimensional environments that span multiple rooms and floors (shown in Figure LABEL:fig:main_splash), these tasks require the seamless integration of high-level planning with low-level motion control.
+For example, an agent may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room.
+Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe agent might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
+The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, sequential dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with behavioral cloning approaches [26].
+Unlike fixed single-arm setups, MoMaâs expansive spaces force the agent to contend with a higher degree of variability and uncertainty.
+Consequently, our benchmark focuses on pick-and-place tasks, such as the ones presented in Figure 1, which serve as a representative foundation for addressing the broader demands of MoMa models in complex, real-world environments.
+Crowdsourced Free-Form Natural Language Instructions.
+Understanding language is crucial for robots in human environments, as it is the most natural way for humans to communicate task specifications.
+Thus, âs tasks are language-conditioned, meaning goals are not specified via images, state information, or symbolic representations.
+Language adds complexity, requiring models to interpret and ground linguistic expressions [27] instead of relying on explicit visual or spatial cues.
+To ensure tasks in reflect real-world needs, we collected free-form human-generated instructions via crowdsourcing.
+This provides a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings.
+Unlike system-generated templated commands, âs free-form instructions, seen in Figures LABEL:fig:main_splash & 1, vary in phrasing, detail, and structure, requiring models to generalize across a variety of expressions rather than memorizing fixed patterns.
+This pushes models to interpret and reason about task objectives.
+Additionally, these instructions introduce ambiguities and implicit references further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.
 III-B Simulated Dataset
-The dataset for our simulated tasks consists of 521 trajectories collected across 20 rooms in five distinct environments, designed to ensure diversity in both object types and room layouts.
-We utilized the AI2THOR [20] simulator and specifically employed RoboTHOR [21] environments, which feature multiple interconnected rooms suitable for long-horizon tasks.
-For robot control, we used ManipulaTHOR [22], which includes an arm capable of fine manipulation and provides low-level pose data.
-Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions taken during the tasks.
-To generate realistic and diverse instructions, we leveraged Prolific111prolific.com for crowdsourcing to gather 521 natural language commands from 41 participants, reflecting tasks that people would naturally expect robots to perform in indoor settings.
-Using these commands, we teleoperated the robot in the simulator with a custom keyboard-based teleoperation system we developed, ensuring accuracy and consistency throughout the data collection process.
-After teleoperation, each trajectory was replay-verified to confirm correctness and eliminate errors, ensuring high-quality data integrity.
+The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and room layouts.
+We utilized the AI2THORÂ [28] simulator, specifically the RoboTHORÂ [29] environments, which feature connected rooms suitable for long-horizon tasks.
+For robot control, we used ManipulaTHORÂ [30], which includes an arm capable of fine manipulation and provides low-level pose data.
+Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions.
+To generate realistic and diverse instructions, we leveraged Prolific222prolific.com for crowdsourcing to gather natural language commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings.
+We teleoperated the simulated robot with a custom keyboard-based teleoperation system we developed to execute these commands.
+After teleoperation, each trajectory was replay-verified to confirm correctness, ensuring data integrity.
 III-C Real-World Dataset
-The dataset for our real-world tasks comprises 50 trajectories collected across 10 rooms in three distinct environments.
+The dataset for our real-world tasks comprises trajectories collected across rooms in three distinct environments.
 These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities.
 The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building.
-These spaces include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to enhance the generalizability of trained models.
-To handle complex terrain and ensure stability, we selected the Spot [23] quadruped mobile manipulator, equipped with an arm for pick-and-place tasks.
-The quadruped’s design enables it to navigate stairs and maintain balance during task execution, making it ideal for the challenging settings of our chosen environments.
-We teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on 50 natural language commands provided by seven participants.
-The commands, distributed as 20, 15, and 15 across the three environments, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings.
-Task trajectories were collected at 3 Hz to ensure detailed and realistic data.
-IV Evaluation
-IV-A Baseline Models
-To evaluate the and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models.
-Given the growing popularity of imitation learning (IL), particularly BC, we benchmark two state-of-the-art (SOTA) BC models.
-These models provide insight into how data-efficient current BC algorithms are by being tested on , validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success.
-Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP).
-This system, evaluated zero-shot without requiring training, investigates whether non-BC methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
-Finally, we assess the performance of a random agent, wherein actions are uniformly sampled to establish a lower-bound baseline.
-RT-1.
-Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a Transformer architecture to process natural language instructions and RGB image inputs into discretized actions.
-Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
-To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robot’s capabilities.
-For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation to achieve precise control.
-MotionGlot-MoMa MG [24] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [25].
-Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
-We adapt MotionGlot (MG) for MoMa tasks by leveraging ’s demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa.
-MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions.
-This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
-In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [26].
-This unified vocabulary allows for action generation similar to text generation.
-Finally, the instruction-tuning template from MG-MoMa is used for text command-based action generation.
-LIMP.
-Language Instruction grounding for Motion Planning (LIMP) [27] is a neuro-symbolic system for pick-and-place MoMa that integrates Linear Temporal Logic (LTL), foundation models, and novel TAMP algorithms to process natural language instructions into executable TAMP plans.
-By translating natural language instructions into LTL formulas enriched with Composable Referent Descriptors, LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution.
-Unlike learning-based methods, LIMP requires no training, allowing it to be evaluated zero-shot.
-This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
+These include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to challenge model generalizability.
+To handle these complex terrains, we teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants.
+The commands, distributed as , , and across the three environments, respectively, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings.
+The trajectories were collected at Hz.
 | Scene 1 | Scene 2 | Scene 3 | Scene 4 | Scene 5 | Task Gen | |||||||||||||
@@
 | Scenes | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 4 | 4 | 1 | 5 | 5 | 5 |
+IV Evaluation
+IV-A Baseline Models
+To evaluate and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic systems.
+Given the growing popularity of imitation learning, particularly BC, we benchmark two state-of-the-art (SOTA) BC models.
+This provides insight into the data efficiency of current BC algorithms.
+Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP).
+The foundation model modules are trained on large-scale visual and textual data, but the system overall requires no robot demonstration training.
+Thus, we evaluated it zero-shot to assess whether non end-to-end methods can achieve greater data efficiency.
+By comparing its performance to learning-based approaches, this evaluation provides insight into whether such methods can achieve comparable or superior results, offering guidance for future development of data-efficient MoMa methodologies.
+Finally, we assess the performance of a random agent to establish a lower-bound baseline, wherein actions are uniformly sampled.
+RT-1.
+Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a transformer architecture [31] to process natural language instructions and RGB image inputs into discretized actions.
+Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
+To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robotâs capabilities.
+For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation.
+MotionGlot-MoMa.
+MotionGlot (MG) [32] addresses the problem of varying action and observation dimensions in multi-embodiment models [33].
+Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
+We adapt MG for MoMa by leveraging âs demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa.
+MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions.
+This process involves two stages: First, a VQ-VAEÂ [34] is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
+In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2Â [35].
+This unified vocabulary enables action generation akin to text generation, utilizing the instruction-tuning template from MG-MoMa for text command-based action execution.
+LIMP.
+Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans.
+By translating instructions into LTL formulas enriched with Composable Referent Descriptors [36], LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution.
+Unlike learning-based methods, LIMP requires no training on robot demonstrations, allowing it to be evaluated zero-shot.
+This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
 IV-B Experiments.
+To evaluate data efficiency, we assess how well models generalize to novel scenarios using Scene Generalization and Task Generalization, two widely used experimental setups [4, 15].
+These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to new scenarios.
+Scene Generalization reflects real-world deployment challenges where robots must perform tasks in previously unseen environments with different layouts and/or objects, while Task Generalization assesses adaptability to new tasks within seen settings.
+Together, these experiments quantify a modelâs data efficiency by isolating generalization challenges at both the scene and task levels, providing insights into how well models leverage limited training data to achieve robust performance.
 Data Split & Generalization.
-To evaluate generalization capabilities, we conduct experiments using carefully designed data splits for scene generalization and task generalization on the simulation dataset, highlighted in Table II.
-For scene generalization, we employ a 5-fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn.
+To implement these evaluations, we conduct experiments using carefully designed data splits on the simulation dataset, highlighted in Table II.
+For Scene Generalization, we employ a -fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn.
 The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes.
-Specifically, 80% of the trajectories from each of the four remaining scenes are included in the training set, and 20% are allocated to the validation set, following a stratified split to preserve data distribution.
+Specifically, of the trajectories from each of the four remaining scenes are included in the training set, and are allocated to the validation set, following a stratified split to preserve data distribution.
 During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene.
-This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining experimental controls.
+This process ensures that we assess the modelâs ability to generalize to entirely unseen environments while maintaining experimental controls.
 The rotation of held-out scenes allows for a comprehensive evaluation of Scene Generalization across the entire dataset.
 For Task Generalization, the data split is consistent across all five scenes.
-The training set comprises 80% of the trajectories from each scene, the validation set includes 10%, and the test set includes the remaining 10%.
-Unlike scene generalization, the held-out test set in this experiment evaluates the model’s ability to generalize to novel tasks rather than unseen environments.
+The training set comprises of the trajectories from each scene, the validation set includes , and the test set includes the remaining .
+Unlike Scene Generalization, the held-out test set in this experiment evaluates the modelâs ability to generalize to novel tasks rather than unseen environments.
 Hyperparameter tuning is performed, resulting in a single model optimized for Task Generalization on the test set.
 Training.
-To benchmark the data efficiency of modern BC methods, we train the RT-1 and MotionGlot models from scratch without any pretrained parameters on the simulation tasks.
-This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks.
+To benchmark the data efficiency of modern BC methods, we train the RT-1 and MG-MoMa models from scratch without any pretrained parameters on the simulation tasks.
+This approach evaluates how well these models learn from our modest yet realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks.
 Both models are trained using the Scene Generalization and Task Generalization experiments, ensuring their capabilities are evaluated under varying conditions.
 Fine-Tuning.
-To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the simulation tasks.
-Empirical analysis showed that freezing some model layers led to better performance than not freezing any layers, so this approach was used for the final experiment.
-Using this setup, we conducted the task generalization experiment to measure how much pretrained MoMa knowledge impacts RT-1’s ability to adapt efficiently to new MoMa tasks.
-V Discussion
-V-A Results
-Scene Generalization.
-The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively.
-None of the individual scenes achieved success rates exceeding , indicating significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for long-horizon MoMa tasks.
-Both models outperform the random agent, demonstrating their ability to learn from the data and satisfy the minimum performance criterion.
-Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1.
-This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robot’s initial spawn location, reducing the complexity of navigation and manipulation required.
-In contrast, Scene 1 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
-Interestingly, for MG-MoMa, Scenes 1 and 3 were the exact opposite of RT-1, with those being the lowest performers.
-It is unclear why that is the case, but it could be that MG-MoMa fails to recognize nearby objects.
-Similarly, Mg-MoMa’s best performers are RT-1’s worst, Scenes 2 and 5.
-Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that Motionglot was not originally intended for mobile manipulation tasks.
-Task Generalization.
-The success rates, presented in Table IV, for the Task Generalization experiment for RT-1 and Mg-MoMa, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment.
-This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments.
-Additionally, while it is unsurprising that RT-1 outperformed MG-MoMa on unseen demonstrations, it is notable that the two models performed comparably on seen demonstrations, and the margin of RT-1’s outperformance was relatively small.
-This is remarkable, particularly given that MotionGlot was originally designed for cross-embodied motion generation rather than specifically for mobile manipulation.
-We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks.
-Surprisingly, RT-1 achieved a marginally higher success rate on unseen tasks () compared to seen tasks ().
-This indicates that the model’s training did not overfit to the training data, which is a positive outcome.
-One plausible explanation is that the sampler may have inadvertently selected a disproportionate number of especially challenging tasks for the seen evaluation set, rendering them overly difficult for the model despite having been part of the training data.
-Task Generalization with Fine-Tuning.
-Interestingly, the unseen task performance was marginally lower compared to the fully trained model’s performance on unseen tasks, as seen in Table IV.
-This result suggests the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world.
-LIMP.
-The results presented thus far indicate that learning-based models, face significant challenges in addressing the complexities of long-horizon MoMa tasks.
-To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
-The evaluation was conducted in real-world environments using Spot, assessing 45 tasks, including a mix of tasks originally specified for both simulated and real-world environments.
-Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks.
-For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts.
-This enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts.
-As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training, making it relatively more data efficient.
-One key factor contributing to LIMP’s success is its use of pre-generated maps of the environments.
-They provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks.
-Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless autoregressive action decoding used in BC models like RT-1.
-This robustness is critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
-Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive prior knowledge to reason about task requirements and perceptual inputs in a zero-shot manner.
-This capability enables LIMP to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training.
-While LIMP demonstrates considerable promise, it still falls short of perfect task execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks.
-Overall, LIMP represents a promising direction for future system development, blending the strengths of learning-based approaches with those of classical planning.
-V-B Ablations
-Impact of Dataset Size.
-The poor performance observed with learning-based methods is expected to stem from the limited, realistically sized dataset of the benchmark, but this needs to be further examined.
-Will simply scaling the data improve the performance?
-Will these models effectively and efficiently use more data to learn better?
-To address this, we investigate how performance scales when progressively larger subsets of the dataset (, , , and ) are used for training.
-This evaluates whether these models can effectively utilize limited data or require substantially more, directly testing their data efficiency and scalability for long-horizon MoMa tasks.
-The ablation results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset.
-We attribute part of the variation to the inherent complexity of these long-horizon MoMa tasks and the variability in the sampled subsets, where smaller subsets can incidentally include a higher proportion of comparatively easier scenarios.
-Overall, this suggests that simply providing more samples does not guarantee better learning outcomes for these tasks, and it is likely that limited data is not the primary cause of the poor performance.
-Moreover, since the linear increase in data size did not result in proportional improvement, it indicates that the model is unable to efficiently learn from the additional data provided, deeming it data inefficient.
-This raises questions about deeper limitations, such as architectural constraints, warranting further investigation.
+We fine-tune RT-1 and MG-MoMa with pretrained parameters and evaluate their performance on the simulation tasks.
+Empirical analysis showed that for RT-1, fine-tuning with some frozen model layers led to better performance than not freezing any, so this approach was used for the final experiment.
+In contrast, for MG-MoMa, the unfrozen model performed better, leading to the use of fully trainable parameters.
+Using this setup, we conducted the Task Generalization experiment to measure how much pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
+Metric.
+To evaluate model performance, we use a task success rate metric that quantifies the completion of long-horizon MoMa tasks.
+Each task consists of four sequential subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal.
+Each successful subtask contributes one point, with a maximum of four points per task.
+The success rate is calculated by averaging the subtask scores within each task and then averaging across all tasks.
+This metric captures both navigation and manipulation performance, providing insight into a modelâs ability to execute multi-step long-horizon tasks.
 | Scene Gen | ||||||
@@
 | Random | ||||||
+V Discussion
+V-A Results
+Scene Generalization.
+The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively.
+This demonstrates significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for our long-horizon MoMa tasks.
+Both models outperform the random agent, demonstrating at least a minimal capacity to learn something from the data.
+Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1.
+This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robotâs initial spawn location, reducing the complexity of navigation and manipulation required.
+In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
+Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that MG was not originally intended for MoMa tasks.
+Task Generalization.
+Table IV presents the results of our Task Generalization experiments, which assess the modelâs ability to adapt to new tasks in a familiar environments.
+Notably, both RT-1 and MG-MoMa show slightly higher overall performance compared to the Scene Generalization experiments.
+This improvement is likely due to the larger training dataset size (see Table II) and the availability of all scenes during training.
+Moreover, while the performance of the baselines is equal for the seen settings, RT-1 outperforms MG-MoMa by on the unseen environments.
+Meanwhile, the random agent showed no difference in performance between seen and unseen tasks, as expected.
+Task Generalization with Fine-Tuning.
+Interestingly, RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV.
+Similarly, MG-MoMa showed lower performance on seen tasks.
+This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained in the real world.
 | Task Gen | Task Gen FT | |||
@@
 | RT-1 | ||||
-| MG-MoMa | – | – | ||
-| LIMP | – | – | – | |
-| Random | – | – | ||
+| MG-MoMa | ||||
+| LIMP | â | â | â | |
+| Random | â | â | ||
+LIMP.
+The results presented thus far show that learning-based models (BC in particular) face significant challenges in addressing the complexities of âs long-horizon MoMa tasks.
+To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
+The evaluation was conducted in real-world environments using Spot, assessing tasks, including a mix of tasks originally specified for both simulated and real-world environments.
+Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks.
+For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts.
+This enabled us to evaluate LIMPâs versatility across both task types, providing insights into its adaptability to different environmental contexts.
+As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training on robot demonstrations, making it relatively more data efficient.
+One key factor contributing to LIMPâs success is its ability to generate instruction-conditioned semantic maps of the environment.
+These maps provide the robot with precise information about the locations of objects and regions of interest, eliminating the need to search and enabling the interpretation of fine-grained semantics specified in language-conditioned tasks.
+Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless action decoding used in BC models like RT-1.
+This robustness is critical for the tasksâ long-horizon navigation aspect, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
+Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner.
+This capability enables it to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training.
+While it demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks.
+Overall, LIMP represents a promising direction for future MoMa development, blending the strengths of learning-based approaches with those of classical planning.
+V-B Ablations
+Impact of Dataset Size.
+ALOHA Unleashed [37], a bimanual manipulation system trained on K demonstrations, only achieved a success rate of for certain tasks, showing that even with large datasets, high performance is not guaranteed.
+Similarly, Lin et al.
+[38] found that increasing demonstrations generally improves manipulation performance, though with diminishing returns.
+Since our benchmark involves MoMaâwith larger action and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data.
+To investigate this, we pose the following questions: Are the baseline models data inefficient?
+Will scaling the data improve their performance?
+To answer these, we investigate how performance scales when progressively larger subsets of the simulation dataset (, , , and ) are used for training in the Task Generalization experiment.
+This evaluates whether these models can effectively utilize limited data or require more, testing their data efficiency and scalability for long-horizon MoMa tasks.
+The ablation results, summarized in Figure 2, show a positive upward trend in performance for RT-1 (represented by the pink line) as the training data increases, albeit a weak one.
+This trend is evident in its average rate of change (RoC).
+This shows that scaling the data does improve RT-1âs overall performance.
+A similar trend is observed in the language, vision, and multimodal communities [39, 40, 41], making it encouraging to see this pattern emerge in the context of long-horizon MoMa tasks as well.
+Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the fewer the data, the worse it generally performs.
+This raises questions about the source of inefficiency, such as model architecture, explored in the following section.
 Impact of Architecture.
-To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture.
-While MotionGlot was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also relies on a Transformer architecture.
-To explore a more distinct alternative, we developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks.
-The LSTM variant is evaluated under the task generalization setup using subset experiments with , , , and of the dataset, enabling us to analyze how its performance scales with increasing data.
-By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
-As shown in Table V, the LSTM variant exhibits no upward trend when moving from to of the data, indicating it is also data inefficient.
-To compare the trends between the two architectures, we analyze their average rates of change.
-Both exhibit similarly non-positive trends, indicating that the data inefficiency is unlikely to primarily stem from the choice of architecture and seems to span across different architectures, since they are not efficiently making use of more data.
-Furthermore, since RT-1’s performance is somewhat better than the LSTM variant’s, and its RoC is slightly higher, it indicates marginally better data efficiency, further supporting the argument that architecture is unlikely the primary cause of RT-1’s data inefficiency observed in the main experiments.
-Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are simply inherently too complex and challenging for learning-based methods.
-Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this work and is left for the community to explore in future work.
+To assess whether data inefficiency stems from the architecture, we evaluate and compare different architectures.
+While MG-MoMa was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also uses a transformer [31].
+To explore distinct alternatives, we replace RT-1âs transformer with alternate architectures while keeping all other components, such as encoders, unchanged to ensure a controlled comparison.
+Specifically, we replace the transformer with an LSTM [42], which we refer to as RL-1.
+Likewise, we incorporate the Mamba architecture [43], leveraging its ability to handle long sequences [43], and call this RM-1.
+To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation as previously conducted.
+By comparing their trends/RoCs and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations or not.
+As shown in Figure 2, both RL-1 and RM-1 exhibit upward trends, indicating that they did not learn more efficiently than RT-1 and that the data inefficiency is spanning across different architectures.
+Identifying the exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore in the future.
 VI Conclusion
-We introduce LAMBDA (), a challenging benchmark consisting of a realistic amount, 571, of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments.
+We introduce LAMBDA (), a challenging benchmark consisting of a realistic amount, , of language-conditioned, long-horizon MoMa tasks, spanning room-to-room and floor-to-floor navigation in indoor environments.
 Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
-To evaluate the benchmark, we tested multiple models to assess their performance and data efficiency on these complex tasks.
-Learning-based models, RT-1 and MotionGlot, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to ’s long-horizon tasks.
-In contrast, the neuro-symbolic system LIMP significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by ’s tasks.
-We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
+To evaluate the benchmark, we tested multiple baseline models to assess their performance and data efficiency on these complex tasks.
+End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks.
+In contrast, a neuro-symbolic system, LIMP, significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by âs tasks.
+We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in current models and guiding the development of more efficient systems for daily indoor tasks.
 Limitations and Future Work.
-While the inclusion of a quadruped like Spot in the real world tasks highlights the advantages of quadrupedal platforms, the lack of simulated quadruped data limits the ability to leverage diverse environments that simulators can provide, such as a broader range of floor-to-floor trajectories.
-Expanding the benchmark to include simulated quadruped tasks would enable greater generalization and provide richer datasets for evaluation.
-Furthermore, extending the benchmark to include tasks beyond pick-and-place, such as those involving more intricate manipulation, would increase its relevance and applicability to a wider range of robotic systems and models.
-Future work could also benchmark reinforcement learning models, which, while not the focus of this study, are known for their sample inefficiency.
-Evaluating their performance on ’s sparse rewards would offer additional insights into their suitability for long-horizon MoMa tasks and broaden the scope of the benchmark.
-| Model | Avg RoC | ||||
-| RT-1 | |||||
-| LSTM |
+Incorporating a quadruped like Spot in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories.
+Expanding the benchmark to include simulated quadruped tasks would enhance generalization and enrich evaluation datasets.
+In the future, the benchmark could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.
+ACKNOWLEDGMENT
+This work is supported by ONR under grant award numbers N00014-22-1-2592 and N00014-23-1-2794, NSF under grant award number CNS-2150184, and with support from Amazon Robotics.
+We thank Aryan Singh, George Chemmala, Ziyi Yang, David Paulius, Ivy He, Lakshita Dodeja, Mingxi Jia, Benned Hedegaard, Thao Nguyen, Tuluhan Akbulut, Chaerin Min, Selena Williams, and George Konidaris.

### v7

@@
 Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation.
-We leverage LAMBDA to benchmark current end-to-end learning methods and a modular neuro-symbolic approaches that combines foundation models with task and motion planning.
-We find that end-to-end methodsâeven when pretrainedâyield lower success rates, while neuro-symbolic methods perform significantly better and require less data.
+We leverage to benchmark current end-to-end learning methods and a modular neuro-symbolic approach that combines foundation models with task and motion planning.
+We find that learning methods, even when pretrained, yield lower success rates, while a neuro-symbolic method performs significantly better and requires less data.
 I INTRODUCTION
-Improving data efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications.
+Improving data-efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications.
 These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions.
-Significant strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3], however existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6].
+Strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3]; however, existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6].
 Collecting such large-scale datasets is resource-intensive, requiring substantial time, labor, and infrastructure.
-This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse tasks, and the need for generalization across environments.
+This data inefficiency arises from the inherent challenges of MoMa tasks: extensive observation, action, and state spaces, as well as diverse tasks, and the need for generalization across environments.
 Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, leaving a significant gap in evaluating progress in this area.
@@
 Many benchmarks lack real-world data [13, 12, 9, 14, 15, 4] or rely on planner-generated trajectories [16, 10, 17], limiting their practicality.
-Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide quadruped data.
+Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide real quadruped data.
 To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities).
@@
 Simulated environments enable diverse testing scenarios and object variability, while real-world data ensures models are prepared for practical deployment.
-We collected this dataset with a quadruped robot [18] due to its optimal form-factor suitable for navigating complex terrain.
+We collected the real data with a quadruped robot due to its optimal form factor, suitable for navigating complex terrain.
 We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks.
 Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the datasetâs ability to support learning.
-A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
-Overall our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
+A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved a success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
+Overall, our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
 (2) High-quality dataset:âs tasks are paired with human-collected robot demonstrations, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
-(3) Comprehensive evaluation:We benchmark several models, including BC and modular neuro-symbolic approaches, providing a baseline for future model development for MoMa tasks.
+(3) Comprehensive evaluation:We benchmark multiple methods, including BC and modular neuro-symbolic approaches, providing a baseline for future MoMa system development.
 | Benchmark | NL | FF NL | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
@@
 Even fewer benchmarks, such as RoboCasa [9] and PARTNR [11], offer natural language instructions in a free-form, unconstrained format rather than predefined templates.
-This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in-the-wild.
-Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17], do not support long-horizon tasks involving navigation across rooms and floors.
+This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in the wild.
+Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17] do not support long-horizon tasks involving navigation across rooms and/or floors.
 These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations.
@@
 The majority of benchmarks in Table I provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models.
-However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11] lack this attribute, limiting their utility for model training and evaluation.
+However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11], lack this attribute, limiting their utility for model training and evaluation.
 M3Bench [10], M2Diffuser [17], and ManiSkill-HAB [16] depend on planner-generated trajectories.
 These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
-Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed quadruped data, which is essential for tasks requiring stability and adaptability.
-Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in .
-Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios.
+Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed real quadruped data, which is essential for tasks requiring stability and adaptability.
+Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as those found in .
+Their robustness allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios.
 The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
-By addressing these limitations, the benchmark provides a comprehensive benchmark to test the data efficiency of MoMa models.
+By addressing these limitations, provides a comprehensive benchmark to test the data efficiency of MoMa models.
 It integrates language-conditioned tasks, long-horizon trajectories spanning multiple rooms and floors, and both simulated and real-world data.
-Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed quadruped data makes it uniquely suited for advancing robotics in complex real-world settings.
+Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed real quadruped data makes it uniquely suited for advancing robotics in complex settings.
 III BENCHMARK
 The LAMBDA () benchmark is designed to evaluate model data efficiency on long-horizon MoMa tasks.
-features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings.
-The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
+features language-conditioned, multi-room, multi-floor pick-and-place demonstrations spanning a range of diverse environments across simulated and real settings.
 III-A Challenges of Tasks
 Data Efficiency in MoMa.
-We define data efficiency as the ability to achieve relatively strong performance with minimal robot trajectory demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every time step.
+We define data efficiency as the ability to achieve relatively strong performance with minimal robot demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every timestep.
 Non-end-to-end systems are still considered within our definition of data efficiency.
-For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstration trajectories, are not excluded on this basis.
-Unlike very large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
+For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstrations, are not excluded on this basis.
+Unlike large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
 While previous tabletop manipulation methods depended on thousands of demonstrations [23], recent models [1, 2, 3] are more data efficient, requiring only tens of demonstrations.
 In a similar direction, recognizing the increased complexity of MoMa, we provide hundreds of demonstrations for model development and benchmarking.
-This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient, avoiding the excessive resource costs of unrealistically large datasets.
+This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient in out-of-distribution generalization, avoiding the excessive resource costs of unrealistically large datasets.
 Long Horizon Tasks in Complex Spaces.
 MoMa tasks introduce significant complexity by involving larger and more intricate action, observation, state, and task spaces compared to tabletop manipulation.
-Operating in dynamic three-dimensional environments that span multiple rooms and floors (shown in Figure LABEL:fig:main_splash), these tasks require the seamless integration of high-level planning with low-level motion control.
-For example, an agent may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room.
-Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe agent might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
-The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, sequential dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with behavioral cloning approaches [26].
-Unlike fixed single-arm setups, MoMaâs expansive spaces force the agent to contend with a higher degree of variability and uncertainty.
-Consequently, our benchmark focuses on pick-and-place tasks, such as the ones presented in Figure 1, which serve as a representative foundation for addressing the broader demands of MoMa models in complex, real-world environments.
+Operating in environments that span multiple rooms and floors (Figure LABEL:fig:main_splash), these tasks require the integration of high-level planning with low-level motion control.
+For example, a robot may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room.
+Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe robot might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
+The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, temporal dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with BC approaches [26].
+Unlike fixed single-arm setups, MoMaâs expansive spaces force the robot to contend with a higher degree of variability and uncertainty.
+Consequently, our benchmark focuses on pick-and-place tasks (e.g., Figure 1), which serve as a representative foundation for addressing the broader demands of MoMa models in complex environments.
 Crowdsourced Free-Form Natural Language Instructions.
-Understanding language is crucial for robots in human environments, as it is the most natural way for humans to communicate task specifications.
+Understanding language is crucial for robots in human environments, as it is the most natural way for humans to specify tasks.
 Thus, âs tasks are language-conditioned, meaning goals are not specified via images, state information, or symbolic representations.
 Language adds complexity, requiring models to interpret and ground linguistic expressions [27] instead of relying on explicit visual or spatial cues.
-To ensure tasks in reflect real-world needs, we collected free-form human-generated instructions via crowdsourcing.
+To ensure tasks in reflect real-world needs, we collected free-form human instructions via crowdsourcing.
 This provides a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings.
-Unlike system-generated templated commands, âs free-form instructions, seen in Figures LABEL:fig:main_splash & 1, vary in phrasing, detail, and structure, requiring models to generalize across a variety of expressions rather than memorizing fixed patterns.
+Unlike system-generated templated commands, âs free-form instructions (e.g., Figures LABEL:fig:main_splash & 1) vary in phrasing, detail, and structure, requiring models to generalize across variations rather than memorizing fixed patterns.
 This pushes models to interpret and reason about task objectives.
-Additionally, these instructions introduce ambiguities and implicit references further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.
+Additionally, these instructions introduce ambiguities and implicit references, further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.
 III-B Simulated Dataset
-The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and room layouts.
+The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and unique room layouts.
+The object configurations within each environment are static.
+Objects and furniture include, but are not limited to: clocks, bats, basketballs, apples, bowls, plates, forks, couches, TVs, chairs, desks, tables, trash cans, beds, shelves, drawers, lamps, etc.
 We utilized the AI2THORÂ [28] simulator, specifically the RoboTHORÂ [29] environments, which feature connected rooms suitable for long-horizon tasks.
@@
 Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions.
-To generate realistic and diverse instructions, we leveraged Prolific222prolific.com for crowdsourcing to gather natural language commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings.
-We teleoperated the simulated robot with a custom keyboard-based teleoperation system we developed to execute these commands.
-After teleoperation, each trajectory was replay-verified to confirm correctness, ensuring data integrity.
+To obtain realistic and diverse instructions, we used Prolific222prolific.com for crowdsourcing to gather commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings.
+We teleoperated the robot with a keyboard-based teleoperation system we developed, to execute these commands.
+Afterward, each trajectory was replay-verified to confirm correctness.
 III-C Real-World Dataset
-The dataset for our real-world tasks comprises trajectories collected across rooms in three distinct environments.
-These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities.
-The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building.
-These include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to challenge model generalizability.
-To handle these complex terrains, we teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants.
-The commands, distributed as , , and across the three environments, respectively, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings.
+The dataset for our real-world tasks comprises trajectories collected across unique rooms in three distinct environments.
+These environments were chosen for their large size, thereby allowing long-horizon tasks, and their status as workplaces where robots can be particularly useful.
+The first is a three-room laboratory, the second is a floor in a university building consisting of five rooms, and the third spans two adjacent floors connected by stairs within the same building, with eight rooms.
+These include kitchens, lobbies, a classroom, and a staircase, offering diverse objects and unique room layouts to challenge model generalizability.
+To handle these complex terrains, we teleoperated the Spot [18] robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants.
+The commands, distributed as , , and across the three environments, respectively, specified room-to-room and floor-to-floor tasks.
 The trajectories were collected at Hz.
 | Scene 1 | Scene 2 | Scene 3 | Scene 4 | Scene 5 | Task Gen | |||||||||||||
+| # Objects | 17 | 27 | 16 | 17 | 18 | - | ||||||||||||
+| # Rooms | 4 | 4 | 3 | 5 | 4 | - | ||||||||||||
 | Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | Train | Val | Test | |
@@
 IV-A Baseline Models
-To evaluate and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic systems.
-Given the growing popularity of imitation learning, particularly BC, we benchmark two state-of-the-art (SOTA) BC models.
+To evaluate the benchmark and establish its initial baselines, we test four MoMa systems.
+Given the growing popularity of BC, we benchmark two state-of-the-art (SOTA) BC models.
 This provides insight into the data efficiency of current BC algorithms.
 Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP).
-The foundation model modules are trained on large-scale visual and textual data, but the system overall requires no robot demonstration training.
-Thus, we evaluated it zero-shot to assess whether non end-to-end methods can achieve greater data efficiency.
+The foundation models are pretrained on large-scale visual and textual data, but the system overall requires no robot demonstration training.
+Thus, we evaluated it zero-shot to assess whether non-end-to-end methods can achieve greater data efficiency.
 By comparing its performance to learning-based approaches, this evaluation provides insight into whether such methods can achieve comparable or superior results, offering guidance for future development of data-efficient MoMa methodologies.
 Finally, we assess the performance of a random agent to establish a lower-bound baseline, wherein actions are uniformly sampled.
-RT-1.
-Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a transformer architecture [31] to process natural language instructions and RGB image inputs into discretized actions.
-Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
-To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robotâs capabilities.
+Robotics Transformer.
+RT-1 [4] is a MoMa BC model.
+It processes natural language instructions and RGB image inputs into discretized actions by leveraging a transformer [31].
+Its status as an open-source, SOTA learning-based MoMa model makes it suitable for benchmarking on as a baseline.
+To adapt RT-1 for the ManipulaTHOR robot, we modified its action head to align with the robotâs capabilities.
 For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation.
@@
 MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions.
-This process involves two stages: First, a VQ-VAEÂ [34] is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
-In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2Â [35].
+This process involves two stages.
+First, a VQ-VAE [34] learns a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions.
+Second, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [35].
 This unified vocabulary enables action generation akin to text generation, utilizing the instruction-tuning template from MG-MoMa for text command-based action execution.
 LIMP.
-Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans.
+Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), a Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans.
 By translating instructions into LTL formulas enriched with Composable Referent Descriptors [36], LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution.
 Unlike learning-based methods, LIMP requires no training on robot demonstrations, allowing it to be evaluated zero-shot.
-This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
+This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient MoMa approaches.
 IV-B Experiments.
-To evaluate data efficiency, we assess how well models generalize to novel scenarios using Scene Generalization and Task Generalization, two widely used experimental setups [4, 15].
-These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to new scenarios.
-Scene Generalization reflects real-world deployment challenges where robots must perform tasks in previously unseen environments with different layouts and/or objects, while Task Generalization assesses adaptability to new tasks within seen settings.
+To evaluate data efficiency, we assess the models on scene generalization and task generalization, two widely used experiment types [4, 15].
+These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to novel scenarios.
+We define Scene Generalization as robots performing tasks in unseen environments with different layouts and/or objects, while Task Generalization as the ability to adapt to new tasks within seen settings.
 Together, these experiments quantify a modelâs data efficiency by isolating generalization challenges at both the scene and task levels, providing insights into how well models leverage limited training data to achieve robust performance.
@@
 Specifically, of the trajectories from each of the four remaining scenes are included in the training set, and are allocated to the validation set, following a stratified split to preserve data distribution.
-During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene.
+During each fold, we train and hyperparameter-tune a separate model specific to the held-out test scene.
 This process ensures that we assess the modelâs ability to generalize to entirely unseen environments while maintaining experimental controls.
@@
 The training set comprises of the trajectories from each scene, the validation set includes , and the test set includes the remaining .
-Unlike Scene Generalization, the held-out test set in this experiment evaluates the modelâs ability to generalize to novel tasks rather than unseen environments.
+Unlike Scene Generalization, the held-out test set in this experiment is fixed and evaluates the modelâs ability to generalize to novel tasks rather than unseen environments.
 Hyperparameter tuning is performed, resulting in a single model optimized for Task Generalization on the test set.
 Training.
-To benchmark the data efficiency of modern BC methods, we train the RT-1 and MG-MoMa models from scratch without any pretrained parameters on the simulation tasks.
-This approach evaluates how well these models learn from our modest yet realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks.
-Both models are trained using the Scene Generalization and Task Generalization experiments, ensuring their capabilities are evaluated under varying conditions.
+We train RT-1 and MG-MoMa without pretrained parameters on the simulation tasks.
+This evaluates how well they learn from our realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks.
+Both models are trained using the Scene Generalization and Task Generalization experiments.
 Fine-Tuning.
-We fine-tune RT-1 and MG-MoMa with pretrained parameters and evaluate their performance on the simulation tasks.
-Empirical analysis showed that for RT-1, fine-tuning with some frozen model layers led to better performance than not freezing any, so this approach was used for the final experiment.
+We fine-tune RT-1 and MG-MoMa with pretrained parameters on the simulation tasks.
+Our empirical analysis showed that for RT-1, fine-tuning with some frozen layers led to better performance than not freezing any, so this approach was used for the final experiment.
 In contrast, for MG-MoMa, the unfrozen model performed better, leading to the use of fully trainable parameters.
-Using this setup, we conducted the Task Generalization experiment to measure how much pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
+This discrepancy may stem from MG-MoMa needing to override task-irrelevant biases from its original non-MoMa pretraining, necessitating updates across all layers.
+In contrast, since RT-1 was already pretrained on MoMa tasks, it may retain useful representations relevant to tasks in certain layers.
+Using this setup, we conducted the Task Generalization experiment to measure the extent to which pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
 Metric.
-To evaluate model performance, we use a task success rate metric that quantifies the completion of long-horizon MoMa tasks.
-Each task consists of four sequential subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal.
+We use a task success rate metric that quantifies the completion of long-horizon MoMa tasks.
+Each task consists of four sequential dependent subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal.
 Each successful subtask contributes one point, with a maximum of four points per task.
 The success rate is calculated by averaging the subtask scores within each task and then averaging across all tasks.
-This metric captures both navigation and manipulation performance, providing insight into a modelâs ability to execute multi-step long-horizon tasks.
+This captures both navigation and manipulation performance, providing insight into a modelâs ability to execute long-horizon tasks.
 | Scene Gen | ||||||
@@
 Scene Generalization.
-The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively.
+The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa, respectively.
 This demonstrates significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for our long-horizon MoMa tasks.
@@
 This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robotâs initial spawn location, reducing the complexity of navigation and manipulation required.
-In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
+In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object.
+Additionally, Scene 2 contained the highest number of objects compared to the other scenes (see Figure II).
+These factors combined significantly increased task complexity.
 Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that MG was not originally intended for MoMa tasks.
 Task Generalization.
-Table IV presents the results of our Task Generalization experiments, which assess the modelâs ability to adapt to new tasks in a familiar environments.
+Table IV presents the results of our Task Generalization experiments.
 Notably, both RT-1 and MG-MoMa show slightly higher overall performance compared to the Scene Generalization experiments.
@@
 Task Generalization with Fine-Tuning.
-Interestingly, RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV.
-Similarly, MG-MoMa showed lower performance on seen tasks.
-This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained in the real world.
+RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV.
+Overall, seen and unseen results were similar, indicating that the models did not overfit.
+However, MG-MoMa exhibited relatively lower performance on seen tasks compared to unseen tasks when fine-tuned.
+This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained on real-world data.
 | Task Gen | Task Gen FT | |||
@@
 LIMP.
-The results presented thus far show that learning-based models (BC in particular) face significant challenges in addressing the complexities of âs long-horizon MoMa tasks.
+The results presented thus far show that learning-based models, BC in particular, face significant challenges in addressing the complexities of âs long-horizon MoMa tasks.
 To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
-The evaluation was conducted in real-world environments using Spot, assessing tasks, including a mix of tasks originally specified for both simulated and real-world environments.
-Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks.
-For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts.
+The evaluation was conducted in real-world environments using Spot, as LIMP was originally designed for real-world execution.
+We utilized tasks, drawn from a mix of simulated and real-world tasks.
+This mixed evaluation was necessary to test its performance on simulation-specified tasks as well.
 This enabled us to evaluate LIMPâs versatility across both task types, providing insights into its adaptability to different environmental contexts.
+For the simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts.
 As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training on robot demonstrations, making it relatively more data efficient.
@@
 This robustness is critical for the tasksâ long-horizon navigation aspect, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
-Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner.
+Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner.
 This capability enables it to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training.
-While it demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks.
+While LIMP demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks.
+In particular, revealed that LIMP exhibits an object-centric bias; it lacks an explicit understanding of spatial constructs such as rooms (e.g., it can navigate to a âcupâ but not to a âkitchenâ).
+It also lacks a notion of the starting location, rendering it unable to interpret instructions that require returning to its point of origin.
+Moreover, due to its reliance on a VLM, its performance is constrained by the VLMâs perceptual limitations, which in this case led to failures in recognizing or locating target objects.
 Overall, LIMP represents a promising direction for future MoMa development, blending the strengths of learning-based approaches with those of classical planning.
@@
 [38] found that increasing demonstrations generally improves manipulation performance, though with diminishing returns.
-Since our benchmark involves MoMaâwith larger action and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data.
+Since our benchmark involves MoMaâwith larger action, state, and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data.
 To investigate this, we pose the following questions: Are the baseline models data inefficient?
@@
 This shows that scaling the data does improve RT-1âs overall performance.
-A similar trend is observed in the language, vision, and multimodal communities [39, 40, 41], making it encouraging to see this pattern emerge in the context of long-horizon MoMa tasks as well.
-Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the fewer the data, the worse it generally performs.
-This raises questions about the source of inefficiency, such as model architecture, explored in the following section.
+This pattern is consistent with what is observed in the language, vision, and multimodal communities [39, 40, 41].
+Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the less data, the worse they generally perform.
+This raises questions about the source of inefficiency, such as model architecture.
 Impact of Architecture.
@@
 To explore distinct alternatives, we replace RT-1âs transformer with alternate architectures while keeping all other components, such as encoders, unchanged to ensure a controlled comparison.
-Specifically, we replace the transformer with an LSTM [42], which we refer to as RL-1.
-Likewise, we incorporate the Mamba architecture [43], leveraging its ability to handle long sequences [43], and call this RM-1.
-To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation as previously conducted.
-By comparing their trends/RoCs and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations or not.
-As shown in Figure 2, both RL-1 and RM-1 exhibit upward trends, indicating that they did not learn more efficiently than RT-1 and that the data inefficiency is spanning across different architectures.
-Identifying the exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore in the future.
+Specifically, we replace the transformer with an LSTM [42], and refer to the model as RL-1.
+Likewise, we incorporate the relatively new Mamba architecture [43], and call this RM-1.
+To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation previously conducted.
+By comparing their trends/RoCs, subsets, and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations.
+As presented in Figure 2, both RT-1 and RL-1 exhibit upward trends with similar RoCs.
+This indicates that RL-1 did not learn more efficiently than RT-1.
+However, RM-1 maintained a relatively constant trend and outperformed RT-1 across all subsets and the full dataset.
+This finding shows that architecture could be a contributing factor in the observed data inefficiency associated with -type tasks.
+Identifying more concrete and exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore.
 VI Conclusion
@@
 Each task is paired with human-collected expert demonstrations from both simulated and real-world settings.
-To evaluate the benchmark, we tested multiple baseline models to assess their performance and data efficiency on these complex tasks.
-End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks.
+To evaluate the benchmark, we tested baseline models to assess their performance and data efficiency on these complex tasks.
+End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks.
 In contrast, a neuro-symbolic system, LIMP, significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by âs tasks.
-We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in current models and guiding the development of more efficient systems for daily indoor tasks.
+We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in models and guiding the development of more efficient MoMa systems.
 Limitations and Future Work.
-Incorporating a quadruped like Spot in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories.
-Expanding the benchmark to include simulated quadruped tasks would enhance generalization and enrich evaluation datasets.
-In the future, the benchmark could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.
+Incorporating a quadruped in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories.
+Expanding the benchmark to include that would enhance generalization and enrich evaluation datasets.
+In the future, could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.
 ACKNOWLEDGMENT